# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Software Development

### v3

@@ 4.2.1. AI-generated code review and compliance
As $RI_{118}$ noted, when code is not generated by the developer, review can be a challenge: “[T]hat’s another big part of it is just going through line by line, but it’s like a more passive mindset, which I find a bit challenging compared to actually writing the code, so it can be harder to actually spot what’s wrong with it.” Accordingly, two interview respondents believed that AI could assist with reviewing AI-generated code. $RI_{149}$ told us that they “ask the GenAI itself to review the code, perhaps using a different model […] [and] ask it to give me gotchas or things to look out for.” $RI_{31}$ tentatively agreed but acknowledged the challenges: “I’ve thought about this […,] and [if] it’s something where the answer is using machine learning to check whether or not the code that you’ve used appears, and would put you at legal risk, […] you’re relying on a large language model to check whether your large language model puts you at legal risk.”
@@
Finding 6: Among respondents, GenAI code review is typically done manually or by using other GenAI to check for code correctness and compliance. Some developers are concerned that the process of conducting a review may itself expose them or their employer to liability.
@@ 4.2.2. Documentation of GenAI usage
One might expect that organizations that permit using GenAI for code development would ensure that developers were aware of processes to document its use so that the organization’s legal risk could be assessed by appropriate personnel. However, of the 445 developers indicating that their organization (partially) permitted the use of GenAI, only 83 (19%) indicated that they were aware of such a process, and of the remaining group, 312 (70%) said that there was no process and another 50 (11%) were unsure if a process existed. $R_{325}$ offered one view of why a process was unnecessary: “I did not use anything that cannot be googled anyway—but we never thought about having to document or question code we found through Google.” Whether this means that many organizations have no process for documenting the use of GenAI or that at least some developers incorrectly believe that no documentation process exists at their organization, the result may be that developers do not have a full understanding of the possible legal risks they or their organizations face (as described in Section 2.1).
Finding 7: Most developers in our study (81%) reported that there was no process for documenting GenAI usage within their organization or that they were unaware of any such process. To the extent that organizations have processes in place for documenting the use of GenAI, they may not be sufficiently educating their employees about those processes.
We further asked the 83 respondents who reported that their companies had processes to document their AI usage about what information was collected during the process. The majority of responses (49, 59%) indicated that documentation consisted of noting the use of an AI tool, with 36 respondents (43%) indicating that the prompt was also documented. In 27 cases (33%), the entire interaction with the code generation tool was documented. Other documentation strategies included keeping system/configuration records (3), summarizing the purpose of using AI tooling (2), providing links to chats (1), documenting known risks associated with the use of AI (1), storing the AI agent’s memory (1), and relying on version control (1).
Finding 8: The most common documentation process among respondents was to note the fact of GenAI usage and/or to document the prompt used to acquire the generation. However, nearly a third of respondents (33%) indicated documenting the whole interaction, potentially leading to reproducibility.
@@ 4.2.3. Review of the Terms of Service for AI Tools
The Terms of Service (ToS) for GenAI tools lay out how developers may interact with models by describing permissible and impermissible behaviors. These ToS may contain language that prevents the use of the service for any illegal activity, forbids the use of model output as training data for another competing model, or (for open-weight models) provides stipulations on what is allowed when fine-tuning. The ToS may also include stipulations that data provided as prompts by a user may be used as training data for future models. In short, it is important for both users and model fine-tuners to understand the terms that they are agreeing to before using a model.
Developers’ views on the ToS of GenAI tools varied widely. A little over a third of respondents who said they incorporated AI into their workflow (177, 36%) indicated that they never read terms in full but may skim them looking for relevant or problematic provisions. A similar percentage of respondents (141, 28%) stated that they sometimes read the ToS in full, depending on the tool they are using. An additional 127 respondents (26%) claimed that they never read the ToS for the tools they use. The smallest group, 52 participants (11%), told us that they always read the ToS in full.
@@
Still other respondents indicated that they relied on others to read the documents and alert them to relevant provisions. Respondents who worked within organizations stated that they trusted that their organization’s legal department would take responsibility for compliance. $RI_{39}$ stated, “I personally don’t worry about [licensing issues related to generated code] because the legal department where I work has already authorized [the] use of [the LLMs], and it’s kind of their problem from my perspective.”
@@
The group of respondents who stated that they did read the ToS for tools they used reported focusing on how the tool developer acquires and uses both personal data and training data, at least for some tools. $RI_{516}$ explained, “I always take a look to see what [GenAI tools] do with the data, the risks involved in inadvertently disclosing personal or private information into the data.” $RI_{31}$ expressed similar concerns, “The training set, making sure that the training set is—they haven’t trained it on stolen code. And then the other one is making sure that I can run the whole thing locally and maintain privacy.” Some others, for example, $RI_{190}$, took a more pragmatic approach, stating, “[The major red flags] would have to be something that could bring me or the company I work at into big trouble, and that would [have to] be more trouble than that we probably won’t exist in a year because our product isn’t good enough, so it would have to be pretty big.”
Finding 9: Even though some provisions in an AI tool’s Terms of Service may be important to developers and their work, developers do not always thoroughly read the Terms of Service for the tools they use, choosing instead to skim the document, assume that their organization has read the document, or rely on information and validation from others in the development community.
@@ 4.2.4. Copyright/legal training
5. RQ2: Developers’ perceptions of licensing and copyright issues emerging from the use of Generative AI
RQ2 sought to determine developers’ beliefs and understandings regarding the licensing and copyright issues that arise from using GenAI tools. Here, we explore developers’ thoughts on current litigation and the prospect of regulation, sentiments regarding the use of their own code in model training data, opinions on the ownership of GenAI output and prompts, and the perceived risk of copyright infringement. This research question is answered by Findings 11-24, which are summarized in our replication package (ano, 2024).
@@ 5.1. Perceptions about litigation and regulation
We asked respondents whether they believed that litigation would result in limitations on the use of GenAI for software development. (Figures 7 and 8 show respondents’ answers to nine five-point Likert-scale (Oppenheim, 2000) questions.) Of the 554 respondents who indicated a familiarity with AI tools, 205 (37%) somewhat agreed with the view that current litigation will result in limitations placed on using GenAI models for software development, with 76 (14%) respondents strongly agreeing. An additional 109 (20%) respondents were neutral, and the third largest group of 104 developers (19%) somewhat disagreed. The smallest group, 60 developers (11%), strongly disagreed.
While more respondents believed that litigation would result in limitations on the use of GenAI, developers who believed that litigation would not result in limitations were more vocal about their views in their responses to open-ended follow-up questions. Indeed, many respondents indicated that technology would outpace attempts to regulate it, whether through litigation or otherwise. $R_{98}$ commented, “I don’t see any limitations coming due to ongoing litigation—yes, it may cause issues for OpenAI and StabilityAI, but you can’t put the genie back in the lamp.” $R_{507}$ further elaborated: “Pandora’s box has been opened. Even if the current models were legally limited, good luck proving that certain code has been used as training data, and good luck stopping people from running Llama instances locally.” $R_{54}$ acknowledged the existence of legal issues but believed that “the momentum that’s already been built by OpenAI and all the hype around the tools means there will not be the political capital to get any restrictive regulations in place.”
@@
Finally, two respondents expressed that they thought it would not be technology that changed to adapt to the law but rather the law that would change to adapt to technology. $R_{376}$ commented, “The result […] will probably not be limitations on the technology since open-source is rapidly developing and cannot be genuinely regulated in this fashion. Rather, intellectual property, artistic rights, etc. will need to be recalibrated.” $R_{182}$ offered an aspirational approach, stating that they “don’t think it is wise to approach the question of copyright of AI-generated/assisted code as a matter of whether the generated code is or is not copyrightable based primarily on what has been settled. Rather, we should think about what we want it to be, keeping in mind the benefits copyright is intended to bring to humanity.”
Finding 11: While about half of participants anticipated that litigation would result in limitations on the use of GenAI, a vocal minority suggested that imposing such limitations would be difficult if not impossible, and others urged restraint, given the pace of technological developments.
@@ 5.2. Using one’s code in models’ training data or as output
If regulation of GenAI occurs, whether through legislation or litigation, stakeholders’ perceptions and practices, including software developers, should ideally be considered. Our study, therefore, sought to learn respondents’ views regarding the use of code as training data for generative models.
@@ 5.2.1. Sentiments regarding the inclusion of one’s own code in training data
Of the 554 respondents familiar with GenAI tools, the largest proportion (41%), as seen in Figure 8, indicated that they would be neither pleased nor displeased if the code they developed was included in training data without their permission. However, open-ended responses indicate that the reality is context-dependent. If the work was open-sourced, respondents generally indicated that they would be pleased or even flattered to have their code included in training data as a way in which they could contribute to the greater good. $R_{144}$ expressed this sentiment: “If it’s just a class, or a few elements from my code, I would actually be pleased and proud that my code is having [an] impact on other peoples’ work and on the advancement of science in general.” $R_{67}$ elaborated, saying, “If the code has been used [in training data] with my implicit (e.g., through BSD-licensed code) or explicit (e.g., through my company’s or my permission) then I would be happy to have contributed. Otherwise [my] answer becomes ‘displeased.’ ” (Some respondents, however, expressed concerns about training models on open-source code given the quality of the training data. $R_{196}$ stated, “[M]ost of the code I write is bad or middling, and the lack of intentional design I’ve seen in the code of open-source projects […] really worries me […] [in that] we can just automatically generate unclear and potentially insecure code.”)
@@
Conversely, developers indicated disappointment about the idea of their proprietary code being included in training data. $R_{82}$ said, “I’m perfectly OK if a company trains their model on MIT [licensed] projects that I have on GitHub, but GitHub training their model on [proprietary] code is a bit different.” Similarly, $R_{98}$ indicated that for private projects, “I wouldn’t be happy if the code was used for LLM training mostly because some of these repositories may contain committed secrets,” and for work projects, “the code in these is in private repositories and everything that’s inside is a corporate secret. If that code is used for LLM training, it could harm my employer by exposing secrets and algorithms we use, which would be extremely displeasing.”
@@
Consistent with the traditional justification for copyright law in the United States, $R_{148}$ reasoned that unauthorized use of work as training data for GenAI could disincentivise human creativity, leading to reluctance from developers to contribute to open source: “[This] is more of a case of how to sustain OS[S], than [it is] a case of code being copied.” Indeed, $R_{245}$ said, if developers are not incentivized to create code, “we will end up with a gradual decline in the quality of all intellectual property as current AI systems are not creative enough to produce IP that is significantly different from the training material.” $R_{514}$ offered a similar view: “The real issue to me is that those engines are [parasiting off] websites like [S]tack [O]verflow [and] actually lowering the amount of money they make. If these websites stop […] produc[ing] good answers that [AI tools] can feed [off of], they [the AI tools] all will die.” $R_{604}$ even speculated that “[c]ode generation tools might encourage creative professionals to hoard knowledge, defeating the original purpose of copyright: enabling protected[,] limited-time monetization followed by release to the public domain.”
@@
Finding 12: Developers’ views on having their code included in models’ training data was context-dependent. If the code was open source, many developers were indifferent or even pleased to have their work included; this was not true for proprietary code. Developers also expressed concerns about the quality of code incorporated into the training data and whether increased use of GenAI tools would disincentivize developers from contributing to the OSS community.
@@ 5.2.2. Conditions around use of one’s code in training data
If using another’s code in training data requires permission, the question then becomes what conditions might be imposed, including monetary compensation and attribution.
Eleven respondents utilized the open-ended questions in the survey to provide their thoughts on whether developers should receive monetary compensation for their code being included in training data. Four of these argued that training on open-source software did not necessitate such compensation or, indeed, that the developers of open-source software had waived their right to any compensation. Six identified some conditions under which they viewed it to be appropriate. $R_{159}$, for example, stated, “[I]f my code is free, and the AI model is paid, I feel wronged and deserve compensation.” Lastly, $R_{613}$ believed that developers should always receive some form of compensation: “[T]he developers whose code these big corporations use to train their models must receive some amount of compensation in terms of free credits, [GPU] time[,], etc. on their platform.”
Separate from this group, $R_{433}$ took a broader view: “If monetary compensation were required to train these models, it would put smaller AI development companies out of business and restrict the currently thriving sector to only the Tech Giants. No one else would have the [means] to not only acquire the data but also to do the training.”
@@
Eight respondents stated that it was also important to respect licensing terms, including attribution for their contributions or (for restrictive licenses) reuse under the same terms. As respondents suggested, modes of attribution could be as fine-grained as a citation provided along with the generated code or as coarse as a manifest that included links to all the training data. In the view of respondents, embedding attribution into the AI model would serve two functions: (1) respecting the rights of developers and (2) helping users and organizations ensure that they comply with the licenses of code they use by providing provenance of the output. As to the latter, $R_{456}$ imagined a situation where an organization using GenAI unintentionally reproduced GPL-licensed code verbatim, introducing license compliance issues into the project. At least two respondents felt responsible for avoiding such problems rested with the models. $R_{665}$ noted, “Developers should not bear the burden for […] copyright […] concerns. Tools should not be producing code which could get a developer in trouble.”
Another avenue would be for developers to create or choose new OSS licenses that provide specific conditions for inclusion in training data. Such licenses would enable creators to control how and when their works are used for GenAI. As $R_{587}$ put it: “I think that open source developers that hold strong opinions about the use of their code with AI models need to update their licenses to enforce those desires.”
Finding 14: Respondents generally wanted licensing terms to be complied with by GenAI models, especially when producing code identical to or very similar to work in the training set.

### v5

@@ 4.2.1. AI-generated code review and compliance
As $RI_{118}$ noted, when code is not generated by the developer, review can be a challenge: “[T]hat’s another big part of it is just going through line by line, but it’s like a more passive mindset, which I find a bit challenging compared to actually writing the code, so it can be harder to actually spot what’s wrong with it.” Two interview respondents, accordingly, believed that AI itself could assist with reviewing AI-generated code. $RI_{149}$ told us that they “ask the GenAI itself to review the code, perhaps using a different model […] [and] ask it to give me gotchas or things to look out for.” $RI_{31}$ tentatively agreed but acknowledged the challenges: “I’ve thought about this […,] and [if] it’s something where the answer is using machine learning to check whether or not the code that you’ve used appears, and would put you at legal risk, […] you’re relying on a large language model to check whether your large language model puts you at legal risk.”
@@
Finding 6: Respondents reported that their GenAI code review is typically done manually or by using other GenAI to check for code compliance. Some developers reported concerns that the process of conducting a review may itself expose them or their employer to liability.
@@ 4.2.2. Documentation of GenAI usage
One might expect that organizations that permit the use of GenAI for code development would ensure that developers were aware of any processes to document its use so that the legal risk to the organization could be assessed by appropriate personnel. However, of the 445 developers who indicated that their organization permitted or partially permitted the use of GenAI, only 83 (19%) indicated that they were aware of any documentation process, and of the remaining group, 312 (70%) said that there was no process and another 50 (11%) were uncertain if a process existed. $R_{325}$ offered one view of why a process was unnecessary: “I did not use anything that cannot be [G]oogled anyway—but we never thought about having to document or question code we found through Google.” Whether this means that many organizations have no process for documenting the use of GenAI or that at least some developers incorrectly believe that no documentation process exists at their organization, the result may be that developers do not have a full understanding of the possible legal risks they or their organizations face (as described in Section 2.1).
Finding 7: Most developers in our study (81%) reported that there was no process for documenting GenAI usage within their organization or that they were unaware of any such process. To the extent that organizations have processes in place to document the use of GenAI, they may not be adequately educating their employees about these processes.
We further asked the 83 respondents who reported that their companies had processes to document their AI usage on what information was collected during the process. (Respondents could select multiple options and provide additional items; the results are presented in Figure 5.) The majority of the responses (49, 59%) indicated that documentation consisted of noting the use of an AI tool, and 36 respondents (43%) indicated that the prompt was also documented. In 27 cases (33%), the entire interaction with the code generation tool was documented. Other documentation strategies included keeping system/configuration records (3), summarizing the purpose of using AI tooling (2), providing links to chats (1), documenting known risks associated with the use of AI (1), storing the AI agent’s memory (1), and relying on version control (1).
Finding 8: The most common documentation process respondents reported was to note the fact of GenAI usage and/or to document the prompt used to acquire the generation. However, almost a third of the respondents whose companies had documentation processes (33%) indicated that they documented the entire interaction, potentially leading to reproducibility.
@@ 4.2.3. Review of the Terms of Service for AI Tools
The Terms of Service (ToS) for GenAI tools convey how developers can interact with models by describing permissible and impermissible behaviors. These ToS may contain language that prevents the use of the service for any illegal activity, forbids the use of model output as training data for another competing model, or (for open-weight models) provides guidance on what is allowed when fine-tuning. The ToS may also provide notice that data provided as user prompts may be used as training data for future models. In short, it is important for both users and model fine-tuners to understand the terms that they are agreeing to before using a model.
Developers’ views on the ToS of GenAI tools varied widely, as reflected in Figure 6. A little over a third of respondents who said they incorporated AI into their workflow (177, 36%) indicated that they never read terms in full but might skim them looking for relevant or problematic provisions. A similar percentage of respondents (141, 28%) stated that they sometimes read the ToS in full, depending on the tool they are using. An additional 127 respondents (26%) claimed that they never read the ToS for the tools they use. The smallest group, 52 participants (11%), told us that they always read the ToS in full.
@@
Other respondents indicated that they relied on others to read the documents and alert them to relevant provisions. Respondents who worked within organizations stated that they trusted that their organization’s legal department would take responsibility for compliance. $RI_{39}$ stated, “I personally don’t worry about [licensing issues related to generated code] because the legal department where I work has already authorized [the] use of [the LLMs], and it’s kind of their problem from my perspective.”
@@
The group of respondents who stated that they did read the ToS for the tools they used reported focusing on how the tool developer acquires and uses both personal data and training data, at least for some tools. $RI_{516}$ explained, “I always take a look to see what [GenAI tools] do with the data, the risks involved in inadvertently disclosing personal or private information into the data.” $RI_{31}$ expressed similar concerns, “The training set, making sure that the training set is—they haven’t trained it on stolen code. And then the other one is making sure that I can run the whole thing locally and maintain privacy.” Some others, for example, $RI_{190}$, took a more pragmatic approach, stating, “[The major red flags] would have to be something that could bring me or the company I work at into big trouble, and that would [have to] be more trouble than that we probably won’t exist in a year because our product isn’t good enough, so it would have to be pretty big.”
Finding 9: Even though some provisions in an AI tool’s Terms of Service may be important to developers and their work, developers reported not always thoroughly reading the Terms of Service for the tools they use, choosing instead to skim the document, assume that others in their organization have read the document, or rely on information and validation from others in the development community.
@@ 4.2.4. Copyright/legal training
5. RQ2: Developer perceptions of licensing and copyright issues emerging from the use of Generative AI
RQ2 sought to determine developers’ beliefs and understandings regarding the licensing and copyright issues that arise from using GenAI tools. Here, we explore developers’ thoughts on current litigation and the prospect of regulation, sentiments regarding the use of their own code in model training data, opinions on the ownership of GenAI output and prompts, and the perceived risk of copyright infringement. This research question is answered by Findings 11-24, which are summarized in our replication package (ano, 2025).
@@ 5.1. Perceptions about litigation and regulation
We asked respondents whether they believed that litigation would result in limitations on the use of GenAI for software development. (Figures 7 and 8 show respondents’ answers to nine five-point Likert-scale (Oppenheim, 2000) questions.) Of the 554 respondents who indicated familiarity with AI tools, 205 (37%) somewhat agreed with the view that current litigation efforts will result in limitations placed on the use of GenAI models for software development, with 76 (14%) respondents strongly agreeing. An additional 109 (20%) respondents were neutral, and the third largest group of 104 developers (19%) somewhat disagreed. The smallest group, 60 developers (11%), strongly disagreed.
Although more respondents believed that litigation would result in limitations on GenAI use, developers who believed that litigation would not result in limitations were more vocal about their views in their responses to open-ended follow-up questions. Many such respondents indicated that technology would outpace attempts to regulate it, either through litigation or otherwise. $R_{98}$ commented, “I don’t see any limitations coming due to ongoing litigation—yes, it may cause issues for OpenAI and StabilityAI, but you can’t put the genie back in the lamp.” $R_{507}$ further elaborated: “Pandora’s box has been opened. Even if the current models were legally limited, good luck proving that certain code has been used as training data, and good luck stopping people from running Llama instances locally.” $R_{54}$ acknowledged the existence of legal issues but believed that “the momentum that’s already been built by OpenAI and all the hype around the tools means there will not be the political capital to get any restrictive regulations in place.”
@@
Finally, two respondents expressed that they thought it would not be technology that changed to adapt to the law but rather the law that would change to adapt to technology. $R_{376}$ commented, “The result […] will probably not be limitations on the technology since open-source is rapidly developing and cannot be genuinely regulated in this fashion. Rather, intellectual property, artistic rights, etc. will need to be recalibrated.” $R_{182}$ offered an aspirational approach, stating that they “don’t think it is wise to approach the question of copyright of AI-generated/assisted code as a matter of whether the generated code is or is not copyrightable based primarily on what has been settled. Rather, we should think about what we want it to be, keeping in mind the benefits [that] copyright is intended to bring to humanity.”
Finding 11: While about half of the participants anticipated that litigation would result in limitations on GenAI use, a vocal minority suggested that imposing such limitations would be difficult, if not impossible, and others urged restraint, given the pace of technological developments.
@@ 5.2. Using one’s code in models’ training data or as output
If regulation of GenAI occurs, whether through legislation or litigation, the perceptions and practices of stakeholders, including software developers, should be considered. Our study therefore sought to learn respondents’ views on the use of code as training data for generative models.
@@ 5.2.1. Sentiments regarding the inclusion of one’s own code in training data
Of the 554 respondents familiar with GenAI tools, the largest proportion (41%), as seen in Figure 8, indicated that they would be neither pleased nor displeased if the code they developed was included in training data without their permission. However, open-ended responses indicated that the reality is context-dependent. If the work was open source, respondents generally indicated that they would be pleased or even flattered to have their code included in training data as a way in which they could contribute to the greater good. $R_{144}$ expressed this sentiment: “If it’s just a class, or a few elements from my code, I would actually be pleased and proud that my code is having [an] impact on other peoples’ work and on the advancement of science in general.” $R_{67}$ elaborated, saying, “If the code has been used [in training data] with my implicit (e.g., through BSD-licensed code) or explicit (e.g., through my company’s or my permission) then I would be happy to have contributed. Otherwise, [my] answer becomes ‘displeased.’ ” (Some respondents, however, expressed concerns about training models on open-source code given the quality of the training data. $R_{196}$ stated, “[M]ost of the code I write is bad or middling, and the lack of intentional design I’ve seen in the code of open-source projects […] really worries me […] [in that] we can just automatically generate unclear and potentially insecure code.”)
@@
In contrast, developers indicated disappointment about the idea of their proprietary code being included in training data. $R_{82}$ said, “I’m perfectly OK if a company trains their model on MIT [licensed] projects that I have on GitHub, but GitHub training their model on [proprietary] code is a bit different.” Similarly, $R_{98}$ indicated that for private projects, “I wouldn’t be happy if the code was used for LLM training mostly because some of these repositories may contain committed secrets,” and for work projects, “the code in these is in private repositories and everything that’s inside is a corporate secret. If that code is used for LLM training, it could harm my employer by exposing secrets and algorithms we use, which would be extremely displeasing.”
@@
Consistent with the traditional justification for copyright law in the United States, $R_{148}$ reasoned that unauthorized use of work as training data for GenAI could disincentivize human creativity, leading to the reluctance of developers to contribute to open source: “[This] is more of a case of how to sustain OS[S], than [it is] a case of code being copied.” Indeed, $R_{245}$ said, if developers are not incentivized to create code, “we will end up with a gradual decline in the quality of all intellectual property as current AI systems are not creative enough to produce IP that is significantly different from the training material.” $R_{514}$ offered a similar view: “The real issue to me is that those engines are [parasiting off] websites like [S]tack [O]verflow [and] actually lowering the amount of money they make. If these websites stop […] produc[ing] good answers that [AI tools] can feed [off of], they [the AI tools] all will die.” $R_{604}$ even speculated that “[c]ode generation tools might encourage creative professionals to hoard knowledge, defeating the original purpose of copyright: enabling protected[,] limited-time monetization followed by release to the public domain.”
@@
Finding 12: Developers’ views on having their code included in models’ training data were context-dependent. If the code was open source, many developers were indifferent or even pleased to have their work included; this was not true for proprietary code. Developers also expressed concerns about the quality of the code incorporated in the training data and whether increased use of GenAI tools would disincentivize developers from contributing to the OSS community.
@@ 5.2.2. Conditions around use of one’s code in training data
If using another developer’s code in training data requires permission, the question then becomes what conditions might be imposed, including monetary compensation and attribution.
Eleven respondents used open-ended questions in the survey to provide their thoughts on whether developers should receive monetary compensation for their code being included in training data. Four of these argued that training on open-source software did not necessitate such compensation or, indeed, that the developers of open-source software had waived their right to any compensation. Six identified some conditions under which they considered it appropriate. $R_{159}$, for example, stated, “[I]f my code is free, and the AI model is paid, I feel wronged and deserve compensation.” $R_{613}$ believed that developers should always receive some form of compensation: “[T]he developers whose code these big corporations use to train their models must receive some amount of compensation in terms of free credits, [GPU] time[,] etc. on their platform.”
$R_{433}$ took a broader view of the impact of compensation: “If monetary compensation were required to train these models, it would put smaller AI development companies out of business and restrict the currently thriving sector to only the Tech Giants. No one else would have the [means] to not only acquire the data but also to do the training.”
@@
Eight respondents stated that it was also important to respect licensing terms, including attribution for their contributions or (for restrictive licenses) reuse under the same terms. As respondents suggested, the modes of attribution could be as fine-grained as a citation provided along with the generated code or as coarse as a manifest that included links to all the training data. In the view of respondents, embedding attribution into the AI model would serve two functions: (1) respecting the rights of developers and (2) helping users and organizations ensure that they comply with the licenses of code they use by providing provenance of the output. In terms of the latter, $R_{456}$ imagined a situation in which an organization using GenAI unintentionally reproduced GPL-licensed code verbatim, introducing license compliance issues into the project. At least two respondents felt that the models should be responsible for avoiding such problems. $R_{665}$ noted, “Developers should not bear the burden for […] copyright […] concerns. Tools should not be producing code which could get a developer in trouble.”
Another avenue would be for developers to create or choose new OSS licenses that provide specific conditions for inclusion in training data. Such licenses would enable creators to control how and when their works are used for GenAI. As $R_{587}$ put it: “I think that open source developers that hold strong opinions about the use of their code with AI models need to update their licenses to enforce those desires.” $R_{429}$, by contrast, conveyed that developers’ preferences should be honored regardless of the existence of prohibitions in the license: “[I]f an individual or institution wishes to have their data removed from a training dataset, they should have the right to request its removal.” (The response did not suggest how such removal might be accomplished.)
Finding 14: Some respondents generally wanted licensing terms to be complied with by GenAI models, especially when producing code identical to or very similar to work in the training set.